If n=300 and X=120, construct a 90% confidence interval estimate for the population proportion. Round to four decimal places as needed
Answer #1
sample success x = 120
sample size          n= 300
sample proportion p̂ =x/n=120/300= 0.400
std error se= √(p*(1-p)/n) =√(0.4*0.6/300) = 0.0283
for 90 % CI value of z= 1.645
margin of error E=z*std error   = 0.0465
lower bound=p̂ -E                       = 0.3535
Upper bound=p̂ +E                     = 0.4465
from above 90% confidence interval for population proportion =(0.3535 , 0.4465)